OpenStudy (anonymous):

Rewrite the expression as a trigonometric function of a single angle measure.

OpenStudy (anonymous):

\[\frac{ \tan4 \theta + \tan2 \theta }{ 1-\tan4\theta \tan2\theta } \]

OpenStudy (anonymous):

without knowing the answer exactly, i am going to guess that it is \(\tan(6\theta)\)

OpenStudy (anonymous):

a. tan6 theta b. cot6 theta c. tan4 theta d. cot4 theta

OpenStudy (anonymous):

and here is the reason for my guess. look up the "addition angle formula" for \(\tan(\alpha +\beta)\)

OpenStudy (anonymous):

if it is \[\tan(\alpha +\beta)=\frac{\tan(\alpha)+\tan(\beta)}{1-\tan(\alpha)\tan(\beta)}\] then imagine \(\alpha =4\theta\) and \(\beta=2\theta\) and you get it right away

OpenStudy (anonymous):

if it is not , then i am wrong.

OpenStudy (anonymous):

ok it is correct. i just checked

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw